A train from station A with uniform acceleration '$' for some distance and then goes with uniform retardation '#' for some more distance to come to rest at station B.the distance between station A and B is 4km and the train takes 4min to complete this journey. if '$' and '#' are in km(min)^-2 then

a)(1/$)+(1/#)=2
b)(1/$)+(1/#)=4
c)(1/$)+(1/#)=(1/2)

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

let the acceleration time duration be t1, and retardation time be t2

v=u+at

$*t1=#*t2, t1+t2=4min,

s=ut +(1/2)at2

(1/2)($*t12 + #t22)=4km